Main INDEX
Monthly INDEX
PREV
NEXT
User name X. Zheng
Log entry time 00:01:52 on June21,2004
Entry number 126459
keyword=swing shift summary
16:00 recovering dipoles; left dipole with the regulator on; right dipole with regulator off (ramp up the current by hand).
16:20 While both dipoles are still recovering, we asked for 30uA beam
to measure the background.
16:40 Left dipole current has been sitting at 944.35A for 15 minutes. We turn off the regulator and change the current setting from 1055.40 to 950.0A, hoping to wake it up. But the current starts to drop. We turn on the regulator again, the current continues dropping for a few minutes, then starts climbing up. This time it came back to the set current (whatever that corresponds to 3.009 GeV/c) normally.
16:50 Left dipole is back! but the beam is off;
17:00 turned on the beam asym feedback, it was off for some reason;
17:12 beam is back, but there is problem with fast feedback;
18:36 right dipole reached 1079A, turn on regulator;
19:05 right dipole reached the set point (3.005 GeV/c); start production on both arms;
19:20 zero Hall C charge asymmery, see halog 126430 and 126432;
21:05 problem with happex feedback monitor, red box "no update in over 2 minutes"; see entry 126442;
21:41-22:20 position scan, see entry 126452;
22:20
22:40 Lisa moved BHWP out, see 126455 for details;
23:30 start ramping up left septa; ask MCC to turn off beam to help reducing the right septa heating.
End of shift: still waiting for right septa to cool down.
Charge accumulated during this shift (not including the position scan and background study):
0.11 C with left arm only;
0.18 C with both arms.
run summary:
run# start stop   comments
2310 16:24 18:51 background study with both HRS dipoles off, may have problem with beam fast feedback, last 20 minutes no beam;
2311 17:57 18:51 left arm production, right arm background study,
2312 18:53 19:03 same as 2311
2313 19:06 20:00 full production run
2314 20:02 20:57 full prod
2315-2319: position scan (partial)